# Introducing Perturb-ability Score (PS) to Enhance Robustness Against Evasion Adversarial Attacks on ML-NIDS Author’s draft for soliciting feedback: 5 Nov. 2024

### v2

@@ 5 Results
This section outlines our evaluation plan and details our evaluation scenarios, and presents and discusses the results of our experiments.
@@ 5.2 Classification of the Features Using PS (RQ1,2,3)
TABLE IV: The performance of an ANN/Random Forest (RF)/SVM/CNN-based NIDS
@@
Table I presents the distribution of features from the two utilized datasets (UNSW-NB15 and the improved CSE-CIC-IDS2018) across different perturb-ability classes, based on our proposed perturb-ability Score (PS). Features classified as low perturb-ability (green) have a PS of 0, indicating robustness to adversarial manipulation. High perturb-ability features (red) are characterized by a PS greater than or equal to 0.87, marking them as highly perturb-able. We selected the threshold of 0.87 to define highly perturb-able features, as discussed earlier. This threshold is derived from the geometric mean of the five PS components, ensuring that at most one PS field has a value as low as 0.5. In our PS evaluation criteria, a score of 0.5 signifies a condition that tends to reduce the overall PS score, without entirely nullifying it. Using the 0.87 threshold ensures that a feature with at most one such condition is still classified within the high perturb-ability class, thereby reinforcing the rigor of our feature selection criteria.
The remaining features, which are neither in the high nor low perturb-ability classes, fall into the medium perturb-ability class (yellow). Table I also shows the percentages of these features across the perturb-ability classes, as defined by our PS. Approximately 35% of the features in both datasets belong to the high perturb-ability class (red).
Fig. 11 depicts the histogram of PS values for each dataset, with the Y-axis representing the number of features and the X-axis representing the PS score. The distribution of PS scores is illustrated, along with the cutoffs between the classes.
Table II provides a detailed classification of the UNSW-NB15 dataset features according to their perturb-ability scores. Features such as ‘dstip‘ and ‘service‘ are categorized as low perturb-ability (green), implying their robustness against adversarial manipulations. In contrast, features like ‘dur‘ and ‘stime‘ are identified as high perturb-ability (red), indicating their susceptibility to adversarial attacks. Table III presents the feature classification for the improved CSE-CIC-IDS2018 dataset. Medium perturb-ability features (yellow), such as ‘Flow Bytes/s‘, occupy an intermediate position between these extremes. It is important to note that low perturb-ability (green) features do not imply that an attacker cannot modify them. Instead, modifying these features may disrupt the network or interfere with the malicious functionality of an attack, or the attacker may have limited or no access to them. Tables VI, VII, and VIII in Appendix A show the definitions of the features in the used datasets.
@@ 5.3 Performance of ML-NIDS using various combinations of features based on their perturb-ability (RQ4)
Table IV presents the performance of ANN, Random Forest, SVM, and CNN as classifiers in NIDS. The evaluation is conducted on two datasets, with results shown for various feature combinations based on their perturb-ability. The table is structured into three sections: (a) performance with low perturb-ability features (green), (b) performance with low and medium perturb-ability features (green + yellow), and (c) performance with all features (green + yellow + red).
For both datasets, the performance remains consistent across all feature combinations, which is highly promising as it suggests that selecting only low perturb-ability (green) features can still yield excellent model performance. The improved CSE-CIC-IDS2018 dataset demonstrated outstanding performance across all models, metrics, and feature sets.
Similarly, the UNSW-NB15 dataset performed very well across all models, metrics, and selected features, although precision was slightly lower (but still above 0.912, which is excellent) compared to the other dataset. However, this marginal difference in precision was unrelated to our perturb-ability-based feature selection. Precision remained consistent across all selected feature sets. Additionally, we ranked the features using Random Forest to select the most important ones, achieving similar precision results. We also experimented with various balancing techniques, such as SMOTE [32], but observed no significant difference in precision. This consistency further confirms that the precision difference is independent of our perturb-ability-based feature selection.
In a nutshell, Table IV shows that using only low perturb-ability (green) features does not create low-performance models; on the contrary, all of these models have an accuracy and F1 score above 0.9.
It is important to note that in all three cases (a, b, and c) of Table IV, thorough pre-processing was conducted, as discussed in Section section 4.3, and feature engineering was applied to remove spurious or faulty features. For example, IP addresses were not treated as numeric values, as this does not generalize well; instead, basic geolocation information was extracted. Time-based features like timestamps were also dropped, as attacks can happen at any time, making this feature non-generalizable. Identification features (e.g., flow ID) were removed for the same reason. Additionally, we applied feature importance ranking techniques to identify the most influential features in the model, ensuring that these attributes provide meaningful and predictive information about the target variable.
TABLE V: Mapping problem-space evasion adversarial attacks’ traffic morphing techniques to features, the features are colored based on our PS classification.
@@
Vitorino et al.[13] [14] [15] modify various flow attributes such as flow duration, average interarrival time between packets, packet rate (packets per second), average forward packet length, smallest forward segment size, minimum interarrival time between packets, and maximum interarrival time. [13] [14] [15] dur, Sjit, Sload, sbytes, Spkts, Sintpkt smeansz Flow Duration, Fwd IAT Total, Fwd IAT Mean, Fwd IAT Std, Fwd IAT Max,Fwd Packet Length Mean, Fwd IAT Min, Fwd IAT Max, Flow Bytes/s, Flow Packets/s
@@ 5.4 Mapping Traffic Morphing of problem-space attacks to ML-NIDS Features (RQ5)
To validate our proposed PS, we analyzed problem-space evasion adversarial attacks targeting ML-NIDS [11, 12, 13, 14, 15, 16, 17, 18]. Table V presents a detailed mapping of traffic morphing techniques employed in these problem-space adversarial attacks to the corresponding perturbed features within the feature-space of the two datasets utilized in this study. The table outlines the problem-space morphing techniques and maps them to the potentially impacted features. Additionally, it classifies these features according to our perturb-ability Score (PS) system, using a color-coded scheme: low perturb-ability (green), medium perturb-ability (yellow), and high perturb-ability (red).
@@
As shown in Table V, the majority of perturbed or affected features in these problem-space evasion attacks exhibit high perturb-ability scores, classified under the red category. A smaller subset of features shows medium perturb-ability. This highlights the efficacy of our defense mechanism, which strategically eliminates high-perturb-ability features, rendering these types of attacks ineffective. Since the attacker manipulates features that are excluded from the final feature set. In other words, our defense will eliminate these easily perturbed features, which may render these types of attacks ineffective, as the attacker is modifying features that are excluded from the selection, making their changes irrelevant to the feature vector.
Table V highlights a crucial observation: minimal modifications in the problem-space can lead to significant alterations across numerous features in the feature-space. While this may suggest that attackers possess considerable power or strategic advantage, these manipulations often introduce unintended side effects [24], which we term collateral damage.
In practical terms, when an attacker aims to modify a specific feature based on the gradient of the targeted mode, for example, the maximum forward IAT, they might induce a substantial delay between two packets in the flow. However, due to the inherent correlations between features in an ML-based NIDS, this single modification can unintentionally impact a broader set of features. For instance, changing the maximum forward IAT could also affect metrics such as flow duration, packets per second, total forward IAT, and forward IAT mean. This collateral damage does not follow any particular direction of the gradient [1], introducing an element of unpredictability. As a result, these unintentional feature perturbations may negatively influence the success of the attack, potentially undermining the adversary’s objectives.
{mdframed}
Remark: In Problem-space, there is a huge difference between the affected features, and the features the attacker meant to perturb, there are side effect features (collateral damage).
It is crucial to highlight that while most features can be modified in the problem-space, labeling a feature as non-perturb-able means it is difficult to manipulate while complying with NIDS domain constraints.
6 Related Work on Conventional Defenses Against Evasion Attacks on ML-NIDS
There are several approaches to defending ML-NIDS against evasion adversarial attacks. He et al. [24] categorize these defense mechanisms into three primary strategies. The first category is parameter protection techniques, which aim to conceal the ML model’s parameters from adversaries, thereby limiting their ability to exploit model vulnerabilities. An example of this approach is Gradient Masking [33], where the gradient information is obscured to hinder adversarial attacks. The second category includes adversarial sample detection methods, which focus on identifying and filtering out adversarial inputs before they can affect the model’s performance [34]. The third category encompasses robustness optimization techniques, which seek to enhance the resilience of ML models through various strategies. Pre-processing techniques, for instance, aim to eliminate adversarial perturbations before they enter the model. One such technique is Input Randomization [35], where random transformations are applied to input data to obscure potential adversarial modifications. Additionally, adversarial training is a widely recognized method, involving the inclusion of adversarial examples in the training dataset to improve the model’s robustness against such attacks [36, 37].
@@ 7.2 Five Features
Sheatsley et al. [23] argued that limiting the number of features an adversary can manipulate does not completely prevent adversarial attacks. Their findings demonstrated that adversarial samples could achieve a 50% success rate by modifying just five randomly selected features.
We would like to point out that our PS-enabled defense does not aim to reduce the number of features but rather to select features that are non-perturb-able in the problem-space during the feature selection phase. Thus, the fundamental purpose of our suggested defense is to leave attackers with as few perturb-able features as possible (as seen in Table V, the attackers will likely have access to far fewer than five perturb-able features). Moreover, The results by Sheatsley et al. [23] focused on targeting NIDS at the feature layer rather than the problem-space. Real-world network environments are far more complex, and the constraints modeled in their research, such as those related to protocols like TCP and UDP, are a small subset compared to the numerous limitations encountered in the actual problem-space, including side-effect features.
@@ 7.3 Problem-space Evasion Adversarial Attacks are Already Extremely Hard for an Attacker
Creating evasion adversarial attacks against ML-NIDS is an extremely complex task for attackers [1]. ElShehaby et al. [1] developed a comprehensive taxonomy of the challenges of adversarial attacks against ML-NIDS. Their exploration of practicality issues spans various attack settings (White- and Black-Box, Feature- and Problem-Space), highlighting challenges such as attackers’ access to feature vectors, correlations between NIDS features, the impractical assumption that attackers have detailed knowledge of the model and feature extraction techniques or can query ML-NIDS, and the added complexity introduced by the dynamic nature of modern ML-NIDS.
Focusing on problem-space evasion attacks, these typically begin with perturbations in the feature-space. These manipulations are then translated into real-world modifications of network packets, a challenge known as the Inverse Feature-Mapping Problem [8]. This issue creates a substantial obstacle for attackers, often complicating, and in many cases preventing, the straightforward application of gradient-based feature-space attacks to generate problem-space adversarial examples [8]. Additionally, there is no guarantee that an optimal solution in the problem-space will closely resemble the intended adversarial features [24]. Furthermore, early efforts to identify mutation operations relied heavily on trial and error, offering little to no theoretical guidance [24]. He et al. [24] also reported in their survey that the evaluations of adversarial attacks they reviewed predominantly focus on the evasiveness of the attacks rather than their maliciousness. In other words, current adversarial attack research against NIDS does not prioritize maintaining the malicious nature of the flow after problem-space modifications. Moreover, Some problem-space perturbations are counterproductive. For instance, introducing delays between packets to perturb IAT features in DoS (Denial-of-Service) attacks can undermine the attack’s effectiveness, as the reduced packet rate would make the attack more manageable and less disruptive.
@@ 7.4 Feature Reduction as a Defense Against Adversarial Attacks
Previous works have explored feature squeezing or reduction as a defense mechanism against adversarial attacks by narrowing the search space available for attackers [13] [3]. While this approach aims to limit the attack surface, our defense strategy goes beyond mere feature removal. Instead of simply reducing the feature set, we leverage domain-specific constraints to identify and select inherently robust features, using PS. These features are carefully chosen based on their resilience to perturbations, where modifications are likely to compromise the malicious or network functionality of evasion adversarial attacks targeting NIDS. By focusing on features that are not easily perturbed without causing significant operational side effects, our approach offers a more effective defense against problem-space adversarial attacks on NIDS. This makes it much more difficult for attackers to succeed without violating network constraints or compromising their malicious capability.
@@ 7.5 NIDS Datasets
We acknowledge that current NIDS datasets exhibit numerous issues, including poor data diversity, highly dependent features, unclear ground truth, traffic collapse, artificial diversity, and incorrect labels [38]. To address these challenges, we utilized the improved version of CSE-CIC-IDS2018 created by Liu et al. [28]. They identified several significant issues related to feature generation and labeling in the original dataset and addressed these by modifying the CICFlowMeter Tool and adding new features. Furthermore, we implemented thorough data pre-processing techniques to maximize the utility of these datasets and prevent data leakage. However, it is important to note that our aim was not to evaluate NIDS datasets but rather to compare the performance of the best model with access to the entire feature set against the best model utilizing only non-perturb-able features.
@@ 7.6 We are losing information!
Some may argue that utilizing PS to drop certain features could result in the loss of information that may be important for the ML-NIDS. While this concern is valid, it is essential that the application of PS be guided by architects, practitioners, and engineers with substantial domain knowledge. On the other hand, our results might suggest that current literature might be utilizing more information and features than necessary. Table IV presents promising results with only the green features. Another pertinent question is: why rely on features that are easily perturb-able by an outsider?
@@ 7.7 Adversarial Attacks on ML-NIDS Research Direction
During our exploration for this work, we observed common issues in research on adversarial attacks against NIDS. Many studies are overly generous to the attacker, often assuming access to information that is rarely available in real-world settings. Grosse et al.[9] noted that this issue is common across artificial intelligence security applications research generally.
In the NIDS domain, many research studies on evasion adversarial attacks presume full knowledge of the model and its selected feature set, which is not feasible in practice. Moreover, research on black-box evasion adversarial attacks assumes the possibility of querying ML-NIDS, which is likely to be infeasible in real-world scenarios [1]. As established in this paper, problem-space adversarial attacks are more practical than feature-space attacks. However, even for these attacks, existing evaluations tend to focus on their evasiveness rather than their maliciousness [24]. Additionally, research on problem-space adversarial attacks against ML-NIDS has largely overlooked the impact of side effects/collateral damage features [24]. There are also considerable issues within NIDS datasets that must be addressed [38].

### v3

@@ 5.2 Classification of the Features Using PS (RQ1,2,3)
Table 4: The performance of an ANN/Random Forest (RF)/SVM/CNN-based NIDS
@@
Table 1 presents the distribution of features from the two utilized datasets (UNSW-NB15 and the improved CSE-CIC-IDS2018) across different perturb-ability classes, based on our proposed perturb-ability Score (PS). Features classified as low perturb-ability (green) have a PS of 0, indicating robustness to adversarial manipulation. High perturb-ability features (red) are characterized by a PS greater than or equal to 0.87, marking them as highly perturb-able. We selected the threshold of 0.87 to define highly perturb-able features, as discussed earlier. This threshold is derived from the geometric mean of the five PS components, ensuring that at most one PS field has a value as low as 0.5. In our PS evaluation criteria, a score of 0.5 signifies a condition that tends to reduce the overall PS score, without entirely nullifying it. Using the 0.87 threshold ensures that a feature with at most one such condition is still classified within the high perturb-ability class, thereby reinforcing the rigor of our feature selection criteria. The remaining features, which are neither in the high nor low perturb-ability classes, fall into the medium perturb-ability class (yellow). Table 1 also shows the percentages of these features across the perturb-ability classes, as defined by our PS. Approximately 35% of the features in both datasets belong to the high perturb-ability class (red). Fig. 10 depicts the histogram of PS values for each dataset, with the Y-axis representing the number of features and the X-axis representing the PS score. The distribution of PS scores is illustrated, along with the cutoffs between the classes. Table LABEL:fig:NB15PS provides a detailed classification of the UNSW-NB15 dataset features according to their perturb-ability scores. Features such as ‘dstip‘ and ‘service‘ are categorized as low perturb-ability (green), implying their robustness against adversarial manipulations. In contrast, features like ‘dur‘ and ‘stime‘ are identified as high perturb-ability (red), indicating their susceptibility to adversarial attacks. Table 3 presents the feature classification for the improved CSE-CIC-IDS2018 dataset. Medium perturb-ability features (yellow), such as ‘Flow Bytes/s‘, occupy an intermediate position between these extremes. It is important to note that low perturb-ability (green) features do not imply that an attacker cannot modify them. Instead, modifying these features may disrupt the network or interfere with the malicious functionality of an attack, or the attacker may have limited or no access to them. Tables 6, 7, and 8 in Appendix A show the definitions of the features in the used datasets.
@@ 5.3 Performance of ML-NIDS using various combinations of features based on their perturb-ability (RQ4)
Table 4 presents the performance of ANN, Random Forest, SVM, and CNN as classifiers in NIDS. The evaluation is conducted on two datasets, with results shown for various feature combinations based on their perturb-ability. The table is structured into three sections: (a) performance with low perturb-ability features (green), (b) performance with low and medium perturb-ability features (green + yellow), and (c) performance with all features (green + yellow + red).
For both datasets, the performance remains consistent across all feature combinations, which is highly promising as it suggests that selecting only low perturb-ability (green) features can still yield excellent model performance. The improved CSE-CIC-IDS2018 dataset demonstrated outstanding performance across all models, metrics, and feature sets. Similarly, the UNSW-NB15 dataset performed very well across all models, metrics, and selected features, although precision was slightly lower (but still above 0.912, which is excellent) compared to the other dataset. However, this marginal difference in precision was unrelated to our perturb-ability-based feature selection. Precision remained consistent across all selected feature sets. Additionally, we ranked the features using Random Forest to select the most important ones, achieving similar precision results. We also experimented with various balancing techniques, such as SMOTE [13], but observed no significant difference in precision. This consistency further confirms that the precision difference is independent of our perturb-ability-based feature selection. In a nutshell, Table 4 shows that using only low perturb-ability (green) features does not create low-performance models; on the contrary, all of these models have an accuracy and F1 score above 0.9.
It is important to note that in all three cases (a, b, and c) of Table 4, thorough pre-processing was conducted, as discussed in Section section 4.3, and feature engineering was applied to remove spurious or faulty features. For example, IP addresses were not treated as numeric values, as this does not generalize well; instead, basic geolocation information was extracted. Time-based features like timestamps were also dropped, as attacks can happen at any time, making this feature non-generalizable. Identification features (e.g., flow ID) were removed for the same reason. Additionally, we applied feature importance ranking techniques to identify the most influential features in the model, ensuring that these attributes provide meaningful and predictive information about the target variable.
Table 5: Mapping problem-space evasion adversarial attacks’ traffic morphing techniques to features, the features are colored based on our PS classification.
@@
Vitorino et al.[37] [38] [36] modify various flow attributes such as flow duration, average interarrival time between packets, packet rate (packets per second), average forward packet length, smallest forward segment size, minimum interarrival time between packets, and maximum interarrival time. [37] [38] [36] dur, Sjit, Sload, sbytes, Spkts, Sintpkt, smeansz Flow Duration, Fwd IAT Total, Fwd IAT Mean, Fwd IAT Std, Fwd IAT Max,Fwd Packet Length Mean, Fwd IAT Min, Fwd IAT Max, Flow Bytes/s, Flow Packets/s
@@ 5.4 Mapping Traffic Morphing of problem-space attacks to ML-NIDS Features (RQ5)
To validate our proposed PS, we analyzed problem-space evasion adversarial attacks targeting ML-NIDS [21, 22, 37, 38, 36, 39, 24, 8]. Table 5 presents a detailed mapping of traffic morphing techniques employed in these problem-space adversarial attacks to the corresponding perturbed features within the feature-space of the two datasets utilized in this study. The table outlines the problem-space morphing techniques and maps them to the potentially impacted features. Additionally, it classifies these features according to our perturb-ability Score (PS) system, using a color-coded scheme: low perturb-ability (green), medium perturb-ability (yellow), and high perturb-ability (red).
@@
As shown in Table 5, the majority of perturbed or affected features in these problem-space evasion attacks exhibit high perturb-ability scores, classified under the red category. A smaller subset of features shows medium perturb-ability. This highlights the efficacy of our defense mechanism, which strategically eliminates high-perturb-ability features, rendering these types of attacks ineffective. Since the attacker manipulates features that are excluded from the final feature set. In other words, our defense will eliminate these easily perturbed features, which may render these types of attacks ineffective, as the attacker is modifying features that are excluded from the selection, making their changes irrelevant to the feature vector. Table 5 highlights a crucial observation: minimal modifications in the problem-space can lead to significant alterations across numerous features in the feature-space. While this may suggest that attackers possess considerable power or strategic advantage, these manipulations often introduce unintended side effects [23], which we term collateral damage. We coined the term collateral damage features because the term side effect features, commonly used in research, is primarily associated with the unintended consequences of medicines, which are generally beneficial. In contrast, within the context of adversarial attacks, these features are unintentionally perturbed by attackers, and adversarial attacks are generally detrimental. Therefore, we believe that collateral damage features is a more accurate and fitting term. In practical terms, when an attacker aims to modify a specific feature based on the gradient of the targeted mode, for example, the maximum forward IAT, they might induce a substantial delay between two packets in the flow. However, due to the inherent correlations between features in an ML-based NIDS, this single modification can unintentionally impact a broader set of features. For instance, changing the maximum forward IAT could also affect metrics such as flow duration, packets per second, total forward IAT, and forward IAT mean. This collateral damage does not follow any particular direction of the gradient [35], introducing an element of unpredictability. As a result, these unintentional feature perturbations may negatively influence the success of the attack, potentially undermining the adversary’s objectives. It is crucial to highlight that while most features can be modified in the problem-space, labeling a feature as non-perturb-able means it is difficult to manipulate while complying with NIDS domain constraints.
6 Related Work
Defending ML-NIDS against evasion adversarial attacks involves three main strategies, as outlined by He et al. [23]: (1) Parameter Protection: Techniques like Gradient Masking [30] obscure model parameters to limit adversarial exploitation. (2) Adversarial Sample Detection: Methods such as those in [17] identify and filter adversarial inputs. (3) Robustness Optimization: Strategies like adversarial training integrate adversarial examples into the training process to enhance resilience [3, 4].
Beyond these general strategies, researchers have explored defenses that focus on feature manipulation for security applications. For instance, Chen et al. [14] introduced cost-aware training, incorporating domain-specific costs into defense strategies, where manipulating features incurs varying levels of difficulty (e.g., asymmetric costs for increasing or decreasing a feature). Unlike their approach, our method leverages the Perturb-ability Score (PS) as a generalizable metric for evaluating the robustness of NIDS features. PS enables the identification and selection of features that are inherently difficult to perturb, thereby reducing the attack surface without adding overhead to the training or testing phases. Furthermore, our mechanism is completely independent of the ML model and can be applied to any model, unlike the method in [14], which focuses specifically on tree ensemble models.
@@ 7.2 Five Features
Sheatsley et al. [34] argued that limiting the number of features an adversary can manipulate does not completely prevent adversarial attacks. Their findings demonstrated that adversarial samples could achieve a 50% success rate by modifying just five randomly selected features. We would like to point out that our PS-enabled defense does not aim to reduce the number of features but rather to select features that are non-perturb-able in the problem-space during the feature selection phase. Thus, the fundamental purpose of our suggested defense is to leave attackers with as few perturb-able features as possible (as seen in Table 5, the attackers will likely have access to far fewer than five perturb-able features). Moreover, The results by Sheatsley et al. [34] focused on targeting NIDS at the feature layer rather than the problem-space. Real-world network environments are far more complex, and the constraints modeled in their research, such as those related to protocols like TCP and UDP, are a small subset compared to the numerous limitations encountered in the actual problem-space, including side-effect features.
@@ 7.3 Problem-space Evasion Adversarial Attacks are Already Extremely Hard for an Attacker
Evasion adversarial attacks against ML-NIDS are highly complex and impractical for attackers [35]. ElShehaby et al. [35] identified key challenges, including limited attacker access to feature vectors, correlations between NIDS features, restricted knowledge of models and feature extraction, and the dynamic nature of modern ML-NIDS. Problem-space evasion attacks face additional hurdles, such as the Inverse Feature-Mapping Problem, which complicates the translation of feature-space perturbations into realistic network packet modifications [31], often causing problem-space attacks to fail to align with intended adversarial features. Moreover, some problem-space perturbations can sabotage attack efficacy, for example, adding delays as perturbations might reduce the effectiveness of DoS attacks.
@@ 7.4 Feature Reduction as a Defense Against Adversarial Attacks
Feature squeezing and reduction have also been explored as defenses, narrowing the attacker’s search space [37, 25]. However, our approach goes beyond simple feature reduction by applying domain-specific constraints to select resilient features using PS. These features resist perturbations that would either compromise the network’s functionality or render the attack ineffective. By focusing on such features, our defense significantly reduces the effectiveness of problem-space adversarial attacks, making it much harder for attackers to succeed without disrupting network operations or their malicious intent.
@@ 7.5 NIDS Datasets
We recognize that existing NIDS datasets suffer from several limitations, including insufficient data diversity and ambiguities in ground truth labeling [18, 28]. To address these shortcomings, we employed the enhanced version of the CSE-CIC-IDS2018 dataset developed by Liu et al. [28]. Their work identified critical challenges in the original dataset, particularly in feature generation and labeling processes, which they rectified by modifying the CICFlowMeter tool and incorporating additional features.
@@ 7.6 We are losing information!
Some may argue that utilizing PS to drop certain features could result in the loss of information that may be important for the ML-NIDS. While this concern is valid, it is essential that the application of PS be guided by architects, practitioners, and engineers with substantial domain knowledge. On the other hand, our results might suggest that current literature might be utilizing more information and features than necessary. Table 4 presents promising results with only the green features. Another pertinent question is: why rely on features that are easily perturb-able by an outsider?
@@ 7.7 Adversarial Attacks on ML-NIDS Research Direction
During our exploration for this work, we observed common issues in research on adversarial attacks against NIDS. Many studies are overly generous to the attacker, often assuming access to information that is rarely available in real-world settings. Grosse et al.[20] noted that this issue is common across artificial intelligence security applications research generally. In the NIDS domain, many research studies on evasion adversarial attacks presume full knowledge of the model and its selected feature set, which is not feasible in practice. Moreover, research on black-box evasion adversarial attacks assumes the possibility of querying ML-NIDS, which is likely to be infeasible in real-world scenarios [35]. As established in this paper, problem-space adversarial attacks are more practical than feature-space attacks. However, even for these attacks, existing evaluations tend to focus on their evasiveness rather than their maliciousness [23]. Additionally, research on problem-space adversarial attacks against ML-NIDS has largely overlooked the impact of side effects/collateral damage features [23]. There are also considerable issues within NIDS datasets that must be addressed [18].